### v1

@@ 4.4 More potential ringing targets
We conducted a search through the Encyclopaedia of Exoplanetary Systems555https://exoplanet.eu/catalog/ (Schneider et al., 2011) for more ringing candidates with a high eccentricity, short pseudo-synchronous rotation period, large planet-to-star radius fraction and low visual magnitude. Figure 8 shows the observed eccentricity and orbital period of each confirmed exoplanet in the Encyclopaedia with an entry for these two quantities. Some exoplanets are not included in the graph because their eccentricity or orbital period is not given in the catalogue, even if they are available elsewhere in the literature as is the case for TOI-4127 b for example. In order to prioritise planets with well constrained observed parameters, we limit ourselves to transiting systems.
Some planets worth considering in future theoretical or observational studies of ringing are Kepler-1656 b (Brady et al., 2018), TOI-4562 b (Heitzmann et al., 2023) and Kepler-419 b (Dawson et al., 2012), highlighted by the purple stars in Fig. 8. Unfortunately, none of these planets seem as promising a target as HD 80606 b. Kepler-1656 b has a small radius of 0.45 and long pseudo-synchronous period while Kepler-419 b and TOI-4562 b orbit quite large and hot stars. Moreover, all three systems are quite dim with visual magnitudes of 11.67 for Kepler-1656 (Høg et al., 2000), 12.09 for TOI-4562 (Zacharias et al., 2013) and 13.04 for Kepler-419 (Zacharias et al., 2013). Therefore, we recommend to prioritise searching for ringing on HD 80606 b first.
@@ 5 Conclusions
We examined which eccentric transiting exoplanets are likely to display detectable ringing, a series of local maxima in their light curves due to the side that received the heating flash at periastron rotating in and out of view. To do so, we used the 3D global circulation model with full radiative transfer coupling expeRT/MITgcm and adapted it to include non-synchronous rotation and eccentric orbits. We simulated five highly eccentric transiting exoplanets, namely HAT-P-2 b, HD 80606 b, TOI-3362 b, TOI-4127 b and HD 17156 b, with the pseudo-synchronous rotation period of Hut (1981). From these atmospheric models, we generated synthetic light curves in the wavelength range of the Low Resolution Spectroscopy mode of the Mid-InfraRed Instrument aboard JWST.
@@
We found that all five planets in our sample feature an equatorial jet throughout their entire orbit, showing that this dynamical regime is quite robust to changing irradiation due to eccentricity. However, the jets of HD 80606 b and TOI-4127 b, the planets with the longest orbital period in our sample, are only driven during the periastron passage and continue under their inertia while slowly decelerating. The planetary rotation period plays a critical role in the dynamical regime, which is exemplified well by the lack of a consistent equatorial jet when HD 80606 b’s rotation period is decreased to half pseudo-synchronous. The rotation period is also crucial to the ringing phenomenon, as it strongly affects both the amplitude and especially the number of ringing peaks. As a result, an observation of ringing or lack thereof can provide a strong constraint on the rotation period, which is otherwise difficult to get. This would allow one to test the assumption of pseudo-synchronous rotation on highly eccentric planets.
@@
Our synthetic light curves predict local maxima on HD 80606 b and TOI-4127 b, while the other three planets only show some small humps at in the 5 – 12 m wavelength range. TOI-4127 b’s one ringing maximum is not detectable because of its rather dim parent star. For a 9 magnitude star such as HD 80606, JWST’s detection threshold is roughly 25 ppm, which is attained by four ringing peaks on HD 80606 b, making it a promising observation target. However, our models neglect the effects of clouds and obliquity, both of which can obstruct ringing. Therefore, our synthetic light curves likely overestimate the amplitude of the ringing peaks. As a result, it is uncertain how many of the four predicted peaks on HD 80606 b are truly detectable. The timing of these peaks is affected by a complicated interplay between the planetary rotation period, equatorial jet speed, viewing geometry and cooling after the periastron passage. If the ringing lasts longer than an advection timescale however, the spacing between peaks approaches the rotation period, allowing one to extract the rotation period more reliably. An observation run with a baseline stretching seven days after the periastron passage would include all predicted peaks. A five day long observation would still cover three likely detectable peaks, enough to get the spacing between peaks after an advection timescale.
@@
From the study of these five exoplanets, we tentatively deduce what factors facilitate observational ringing. These are a high eccentricity, a low apparent magnitude of the home star, a large planet-to-star radius ratio and a short pseudo-synchronous rotation period. In our sample of five exoplanets the ringing is more prominent on planets with a lower equilibrium temperature and long orbital period, although this goes against previous works, so the role of the orbital period is still unclear. Based on these conditions, we searched for other transiting exoplanets in the Encyclopaedia of Exoplanetary Systems that may display observable ringing, but did not find a candidate as promising as HD 80606 b. Therefore, we conclude that HD 80606 b is currently the best known transiting exoplanet candidate for ringing.
@@ Acknowledgements.
The research presented in this paper was part of MV’s master thesis, supervised by LD and co-supervised by RB, ADS and LC. ADS, LD and LC acknowledge funding from the European Union H2020-MSCA-ITN-2019 under Grant no. 860470 (CHAMELEON). LD further acknowledges support from the KU Leuven IDN/19/028 grant ESCHER. RB acknowledges support from the ’Origins’ investment incentive. The authors would like to thank the unknown referee for their helpful comments. The data presented in this work were collected using the ‘pleiads’ computer cluster of the Institute of Astronomy, KU Leuven, maintained by Maarten Dirickx. For this work we have made use of the GCM post-processing library gcm_toolkit Schneider et al. (2022b).References

### v2

@@ 4.4 More potential ringing targets
We conducted a search through the Encyclopaedia of Exoplanetary Systems555https://exoplanet.eu/catalog/ (Schneider et al., 2011) for more ringing candidates with a high eccentricity, short pseudo-synchronous rotation period, large planet-to-star radius fraction and low visual magnitude. Figure 8 shows the observed eccentricity and orbital period of each confirmed exoplanet in the encyclopaedia with an entry for these two quantities. Some exoplanets are not included in the graph because their eccentricity or orbital period is not given in the catalogue, even if they are available elsewhere in the literature as is the case for TOI-4127 b for example. In order to prioritise planets with well-constrained observed parameters, we limited ourselves to transiting systems.
Some planets worth considering in future theoretical or observational studies of ringing are Kepler-1656 b (Brady et al., 2018), TOI-4562 b (Heitzmann et al., 2023), and Kepler-419 b (Dawson et al., 2012), highlighted by the purple stars in Fig. 8. Unfortunately, none of these planets seem as promising a target as HD 80606 b. Kepler-1656 b has a small radius of 0.45 and long pseudo-synchronous period while Kepler-419 b and TOI-4562 b orbit quite large and hot stars. Moreover, all three systems are rather dim with visual magnitudes of 11.67 for Kepler-1656 (Høg et al., 2000), 12.09 for TOI-4562 (Zacharias et al., 2013), and 13.04 for Kepler-419 (Zacharias et al., 2013). Therefore, we recommend prioritising searches for ringing on HD 80606 b.
@@ 5 Conclusions
We examined which eccentric transiting exoplanets are likely to display detectable ringing, which is a series of detectable local maxima in the light curves due to the side of the planet that received the heating flash at periastron rotating into and out of view. To do so, we used the 3D GCM with full radiative transfer coupling expeRT/MITgcm and adapted it to include non-synchronous rotation and eccentric orbits. We simulated five highly eccentric transiting exoplanets, namely HAT-P-2 b, HD 80606 b, TOI-3362 b, TOI-4127 b, and HD 17156 b, using the pseudo-synchronous rotation period from Hut (1981). From these atmospheric models, we generated synthetic light curves in the wavelength range of the LRS mode of the MIRI aboard JWST.
@@
All five planets in our sample feature an equatorial jet throughout their entire orbit, showing that this dynamical regime is quite robust to changing irradiation due to eccentricity. However, the jets of HD 80606 b and TOI-4127 b, the planets with the longest orbital period in our sample, are only driven during the periastron passage and continue under their inertia while slowly decelerating. The planetary rotation period plays a critical role in the dynamical regime, which is exemplified well by the lack of a consistent equatorial jet when HD 80606 b’s rotation period is decreased to half pseudo-synchronous. The rotation period is also crucial to the ringing phenomenon, as it strongly affects both the amplitude and, especially, the number of ringing peaks. As a result, an observation of ringing or a lack thereof can provide a strong constraint on the rotation period, which is otherwise difficult to get. This would allow one to test the assumption of pseudo-synchronous rotation on highly eccentric planets.
@@
Our synthetic light curves predict local maxima on HD 80606 b and TOI-4127 b, while the other three planets only show some small humps in the 5 – 12 m wavelength range. TOI-4127 b’s one ringing maximum is not detectable because of its rather dim parent star. For a 9 magnitude star such as HD 80606, JWST’s detection threshold is roughly 25 ppm; this is attained by four ringing peaks on HD 80606 b, making it a promising observation target. However, our models neglect the effects of clouds and obliquity, both of which can obstruct ringing. Therefore, our synthetic light curves likely overestimate the amplitude of the ringing peaks. As a result, it is uncertain how many of the four predicted peaks on HD 80606 b are truly detectable. The timing of these peaks is affected by a complicated interplay between the planetary rotation period, equatorial jet speed, viewing geometry, and cooling after the periastron passage. If the ringing lasts longer than an advection timescale, however, the spacing between peaks approaches the rotation period, allowing one to extract the rotation period more reliably. An observation run with a baseline stretching seven days after the periastron passage would include all predicted peaks. A five-day-long observation would still cover three likely detectable peaks, enough to get the spacing between peaks after an advection timescale.
@@
From the study of these five exoplanets, we have tentatively deduced what factors facilitate observational ringing. They are a high eccentricity, a low apparent magnitude of the home star, a large planet-to-star radius ratio, and a short pseudo-synchronous rotation period. In our sample of five exoplanets, the ringing is more prominent on planets with a lower equilibrium temperature and a long orbital period; this contradicts the results of previous works, so the role of the orbital period is still unclear. Based on these conditions, we searched for other transiting exoplanets in the Encyclopaedia of Exoplanetary Systems that may display observable ringing but did not find a candidate as promising as HD 80606 b. Therefore, we conclude that HD 80606 b is currently the best known transiting exoplanet candidate for ringing.
@@ Acknowledgements.
The research presented in this paper was part of MV’s master thesis, supervised by LD and co-supervised by RB, ADS and LC. ADS, LD and LC acknowledge funding from the European Union H2020-MSCA-ITN-2019 under Grant no. 860470 (CHAMELEON). LD further acknowledges support from the KU Leuven IDN/19/028 grant ESCHER. RB acknowledges support from the ’Origins’ investment incentive. The authors would like to thank the unknown referee for their helpful comments. The data presented in this work were collected using the ‘pleiads’ computer cluster of the Institute of Astronomy, KU Leuven, maintained by Maarten Dirickx. In this work, we have made use of the GCM post-processing library gcm_toolkit Schneider et al. (2022b).References